### v2

@@ Abstract
Non-terrestrial network (NTN) is expected to be a critical component of Sixth Generation (6G) networks, providing ubiquitous services and enhancing the system resilience. However, the high-altitude operation and inherent mobility of NTN introduce significant challenges across the development and operations (DevOps) lifecycle. Apart from that, how to achieve artificial intelligence native (AI-Native) capabilities in NTN for intelligent network management and orchestration remains an important challenge. To solve the challenges above, we propose integrating the Open Radio Access Network (ORAN) with NTN as a promising solution, leveraging its principles of disaggregation, openness, virtualization, and embedded intelligence. Despite extensive technical literature on ORAN and NTN, respectively, there is a lack of a holistic view of the integration of ORAN and NTN architectures, particularly in terms of how intelligent ORAN can address the scalability challenge in NTN management. To address this gap, this paper provides a comprehensive and structured overview of an AI-native ORAN-based NTN framework to support dynamic configuration, scalability, and intelligent orchestration. The paper commences with an in-depth review of the existing literature from leading industry and academic institutions, subsequently providing the necessary background knowledge related to ORAN, NTN, and AI-Native for communication. Furthermore, the paper analyzes the unique DevOps challenges for NTN and proposes the orchestrated AI-Native ORAN-based NTN framework, with a detailed discussion on the key technological enablers within the framework. Finally, this paper presents various use cases and outlines the prospective research directions of this study in detail.
@@ 1.1 MOTIVATION
The emerging wireless networks, such as Beyond Fifth Generation (B5G) or Sixth Generation (6G) networks, are envisioned to merge the digital, physical, and human domains using intelligent infrastructures and platforms [B5G-1]. Apart from that, they are also expected to support a variety of applications and use cases that demand high data rates, low latency, and a very large number of users or devices with ubiquitous connectivity [B5G-2, B5G-4].
@@
DRL Deep Reinforcement Learning PDCP Packet Data Convergence Protocol
DT Digital Twin PHY Physical Layer
E2E End-to-End RAN Radio Access Network
eMBB Enhanced Mobile Broadband RF Radio Frequency
FH Fronthaul RIC RAN Intelligent Controller
FL Federated Learning RLC Radio Link Control
FSO Free-Space Optical RRC Radio Resource Control
GEO Geostationary Earth Orbit RT-RIC Real-Time RAN Intelligent Controller
HAPs High-Altitude Platforms SDAP Service Data Adaptation Protocol
ISAC Integrated Sensing and Communication SMO Service Management and Orchestration
@@
However, the deployment of dense 5G networks has revealed several limitations for vendors and mobile network operators (MNOs), including increased network complexity, high capital expenditure (CapEx), and operational expenditures (OpEx). It is noted that energy efficiency has emerged as a critical issue, primarily due to the substantial energy consumption of the Radio Access Network (RAN), which comprises a dense deployment of base stations (BS) and large-scale multiple-input multiple-output (MIMO) systems at each site. More importantly, the coexistence of legacy and non-cloud-native applications, combined with users’ requirements, has further highlighted the need for more intelligent, flexible, and cost-effective solutions in future 6G systems. Specifically, these challenges have motivated significant research into novel network architectures and communication schemes that aim to: (i) Meet the network performance requirements in terms of coverage, data rates, and latency; (ii) Support the envisioned applications such as integrated sensing and communication (ISAC), artificial intelligence (AI) agent-driven interactions and cooperation, and cognitive cities; (iii) Enable the deployment of cost-effective, energy-efficient, scalable, and flexible networks; and (iv) Integrate intelligence and data to optimize the network performance, increase operational efficiency, and improve user experience with a variety of customized services.
@@
The evolution of the RAN towards an open, disaggregated, and intelligent framework, i.e., Open RAN (ORAN), represents a key step in this direction [openRAN-8]. Meanwhile, since B5G and 6G systems become increasingly integrated into important application scenarios, such as smart cities or healthcare, any service disruption can lead to serious socio-economic consequences, especially during emergencies or large-scale failures [khaloopour2024resilience]. To extend coverage and enhance the overall network resilience, the terrestrial network (TN) can be effectively complemented by non-terrestrial network (NTN), which provides complementary and redundant communication links and helps achieve ubiquitous connectivity [alves20256g, 3gppntn]. Nevertheless, the deployment and operation of NTNs, as well as their interoperability with TN, present several challenges. These include the size, weight, and power (SWaP) constraints, excessive propagation delays, the link reliability and mobility management issues, and dynamic resource management. These challenges have motivated increasing interest in the integration of ORAN and NTNs, leveraging ORAN’s inherent features of disaggregation, openness, and AI-Native intelligence to enable more efficient network optimization and innovation [dahrouj2023machine, rihan2023ran]. However, due to the aforementioned constraints and challenges, the adoption of ORAN principles in NTN remains significantly more complicated than in the traditional TN, thus requiring further studies on effective architectural and orchestration solutions.
1.2 REVIEW OF EXISTING RELATED LITERATURE
@@
ORAN-based NTN has recently gained significant attention and research interests, with many research papers advancing the idea through investigations on different aspects. It should be noted that the scope of this subsection does not aim to cover the entirety of the existing works in ORAN for NTN, but rather to specifically review the research papers related to state-of-the-art development and deployment solutions of ORAN-based NTN. Therefore, a detailed content comparison between this review paper and the most relevant research papers is provided in Table 2.
@@
Table 2: Comparison of related works and this review paper on ORAN-based NTN (L: Low, M: Medium, H: High).
@@
[baranda2025architectural] 2025 Survey M L H H L M Analyzes functional splits and RIC placement for ORAN-enabled satellite-based NTNs, highlighting the implementation constraints and interoperability requirements.
[openRAN-NTN-1] 2024 H M H M L M Provides a detailed overview of the utilization of network slicing, AI/ML techniques, and ORAN to address the challenges of NTN systems in both academia and industry.
[mushi2024open] 2024 L L M M H M Investigates the existing ORAN testbeds to support UAVs with controlled mobility, and evaluates AERPAW as a reference platform for future 6G NTN experimentation.
[mahboob2025transforming] 2025 Magazine M L H M L M Proposes an ORAN-based NTN architectures with transparent or regenerative payloads configurations, and outlines the existing challenges in channel, computing and algorithms.
[baena2025space] 2025 M M H H M M Proposes the Space-O-RAN framework, integrating satellite-based NTN and TN based on ORAN and AI-driven solutions, to address dynamic connectivity and scalability issues.
[firouzi20252] 2025 M L M M H H Proposes the Orbital O-RAN (O2-RAN) framework to support multi-layer, multi-segment NTN deployment, and enable intelligent and resilient operations across NTN infrastructures.
[shah2025urllc] 2025 M L L H L L Evaluates different functional split options in ORAN-enabled UAV-based NTNs for URLLC services in terms of computation complexity, latency, and reliability.
[zeydan2025semantic] 2025 M L H M L M Proposes the SEM-NTN framework, integrating semantic communication, ORAN, and NTNs, for real-time, bandwidth-constrained and AI-Native applications.
[moore2025next] 2025 M L M M L M Proposes an AERPAW-based ORAN experimentation platform prototype for wireless UAV research, and analyzes the key challenges and practical solutions for the platform.
[shinde2024ml] 2024 M L L M L L Discusses the options and challenges for deploying ML solutions on ORAN-based TN and NTN scenarios to support centralized and distributed intelligent solutions.
[choi2024spectrum] 2024 L L L M H L Discusses the utilization of ORAN architecture and spectrum marketplaces for facilitating spectrum-sharing strategies between NTNs and TN.
[NTN-challenges-2] 2024 L L M L L M Investigates the ORAN-based NTN architecture in enabling network optimization against various targets, and presents a neutral host 5G network testbed for evaluating the architecture.
[campana2023ran] 2023 L L M L L M Explores the implementation of ORAN-based NTN architecture and investigates the potential trends, such as radio resource management, to enhance NTN system efficiency.
[ORAN_white_paper] 2025 White Paper L M H M M L Presents a detailed overview of the integration of ORAN architecture and 3GPP NTN, emphasizing its importance, current status, challenges, and security considerations.
[pasumarthy2026designing] 2025 Book Chapter L L M M H L Summarizes the 3GPP-based NTN specifications, challenges, and various use cases, showing the benefits for realizing the NTN solutions based on ORAN architecture.
@@
Some of these publications provide a detailed overview of the integration of ORAN and NTNs. For instance, the survey paper [baranda2025architectural] investigated the architectural and functional split strategies and the RAN intelligent controller (RIC) placement options for ORAN-enabled satellite-based NTN systems, and emphasized the implementation constraints and interoperability considerations. Another in-depth survey paper [openRAN-NTN-1] provided a detailed overview of the utilization of network slicing, AI or machine learning (ML) techniques, and ORAN to address the challenges of NTN systems in both academia and industry. Furthermore, the survey paper [mushi2024open] discussed the design requirements of ORAN testbeds for controlled experimentation with UAV-based aerial clients, presented the representative experiments from aerial experimentation and research platform for advanced wireless (AERPAW) related to ORAN, and highlighted the AERPAW platform as a potential future ORAN-based UAV testbed.
Furthermore, several magazine papers focused on addressing the overall architecture related to ORAN-based NTNs [mahboob2025transforming, baena2025space, firouzi20252]. Specifically, the authors in [mahboob2025transforming] proposed an ORAN-empowered NTN architecture with transparent, regenerative-DU, and regenerative payload configurations, and the authors in [baena2025space] proposed the Space-ORAN architecture framework by integrating satellite-based NTN and TN based on ORAN principles and AI-driven solutions. In [firouzi20252], a unified architectural framework named Orbital ORAN (O2-RAN) is proposed to support multi-layer, multi-segment NTN deployment. In addition, other magazine papers have investigated specific aspects of ORAN-enabled NTNs, including ultra reliable low latency communications (URLLC) services [shah2025urllc], semantic communication [zeydan2025semantic], machine learning solution deployment [shinde2024ml], spectrum sharing [choi2024spectrum], and the implementation or testbed designs [moore2025next, NTN-challenges-2, campana2023ran] based on AERPAW. Finally, the white paper [ORAN_white_paper] analyzed the integration of ORAN architecture and 3GPP-based NTN systems in detail, and the book chapter [pasumarthy2026designing] summarized the relevant characteristics for 3GPP NTN systems and highlighted the benefits of integrating NTN with ORAN architecture. Other surveys or magazine papers with limited relevance to the scope of ORAN for NTN can be found in [NTORAN, muro20245g, sun2024advancing, soltani2025intelligent, openRAN-TN-1].
@@
1.3 CONTRIBUTIONS
In this survey article, we provide a comprehensive review of the ORAN solutions for NTNs and analyze the potential of the AI-Native system in enhancing communication performance. Unlike related studies [baranda2025architectural, openRAN-NTN-1, mushi2024open, mahboob2025transforming, baena2025space, firouzi20252, shah2025urllc, choi2024spectrum, zeydan2025semantic, shinde2024ml, moore2025next, NTN-challenges-2, campana2023ran, ORAN_white_paper, pasumarthy2026designing], we analyze the unique DevOps challenges in the NTN lifecycle, propose ORAN-based NTN architectures with key technique enablers, present various use cases, and future directions. The contributions of this paper are summarized as follows:
  • We provide an up-to-date and structured review of the most recent and related studies on the ORAN-based NTN, highlighting their current progress, contributions, and limitations. In addition, we provide a concise background on ORAN, NTN, and AI-Native for communication, and also present key survey and magazine papers for readers interested in exploring these topics further.
  • We identify and analyze the unique challenges across the DevOps lifecycle of NTNs, including SWaP limitations, robust connectivity, platform control and data analytics, and on-orbit test and upgrade. More importantly, we discuss the corresponding ORAN-based solutions by leveraging its inherent features to address these challenges.
  • We then propose a novel orchestrated AI-Native ORAN-based NTN framework to enable dynamic, scalable, and intelligent network operation for future 6G systems. We also highlight its key enabling technologies, including the dynamic wireless fronthaul (FH) split, multi-tier RIC for distributed learning, flexible functional role switch, and cross-domain service management and orchestration.
@@ 1.3 CONTRIBUTIONS
  • We further analyze the future research directions on the integration of the proposed ORAN-based NTN framework with other technologies, including infrastructure sharing, hierarchical ML-based optimization, mobility management, edge AI, cross-domain AI, and digital twin. These directions will further shape the evolution of AI-driven intelligent wireless networks.
1.4 ORGANIZATION
The remainder of this paper is organized as follows. Section 2 provides the detailed background for ORAN, NTN, and AI-Native for communication, and reviews the latest studies related to these topics. Section 3 analyzes the unique challenges of NTN throughout the DevOps lifecycle and discusses corresponding ORAN-based solutions. Section 4 proposes a novel AI-Native ORAN-based NTN framework designed to achieve dynamic and scalable network configuration, along with its key enabling technologies. Section 5 provides three representative use cases empowered by the proposed AI-Native ORAN-based NTN framework. Section 6 outlines future research directions for integrating the ORAN-based NTN framework with other emerging technologies. Finally, Section 7 concludes the paper. An overview of the paper organization is shown in Fig. 1, and the list of key acronyms used is provided in Table 1.

### v3

@@ Abstract
Non-terrestrial network (NTN) is envisioned as a critical component of Sixth Generation (6G) networks by enabling ubiquitous services and enhancing network resilience. However, the inherent mobility and high-altitude operation of NTN pose significant challenges throughout the development and operations (DevOps) lifecycle. To address these challenges, integrating NTNs with the Open Radio Access Network (ORAN) is a promising approach, since ORAN can offer disaggregation, openness, virtualization, and embedded intelligence. Despite extensive literature on ORAN and NTN, a holistic view of ORAN-based NTN frameworks is still lacking, particularly regarding how ORAN can effectively address the existing challenges of NTN. Furthermore, although artificial intelligence native (AI-Native) capabilities have the potential to enhance intelligence network control and optimization, their practical realization in NTNs has not yet been sufficiently investigated. Therefore, in this paper, we provide a comprehensive and structured overview of AI-Native ORAN for NTN. This paper commences with an in-depth review of the existing literature and subsequently introduces the necessary background about ORAN, NTN, and AI-Native for communication. After analyzing the DevOps challenges for NTN, we propose the orchestrated AI-Native ORAN-based NTN framework and discuss its key technological enablers. Finally, we present the representative use cases and outline the prospective future research directions of this study.
@@ 1-A MOTIVATION
The emerging wireless networks, such as Beyond Fifth Generation (B5G) or Sixth Generation (6G) networks, are envisioned to merge the digital, physical, and human domains using intelligent infrastructures and platforms [126]. Moreover, wireless networks are also expected to support a variety of applications and use cases that demand high data rates, low latency, and a very large number of users or devices with ubiquitous connectivity [98, 29].
@@
DRL Deep Reinforcement Learning ORAN Open Radio Access Network
DT Digital Twin PDCP Packet Data Convergence Protocol
E2E End-to-End PHY Physical Layer
eMBB Enhanced Mobile Broadband RAN Radio Access Network
FH Fronthaul RF Radio Frequency
FL Federated Learning RIC RAN Intelligent Controller
FSO Free-Space Optical RLC Radio Link Control
GEO Geostationary Earth Orbit RRC Radio Resource Control
HAPs High-Altitude Platforms RT-RIC Real-Time RAN Intelligent Controller
ISAC Integrated Sensing and Communication SDAP Service Data Adaptation Protocol
LAE Low-Altitude Economy SMO Service Management and Orchestration
@@
However, the deployment of dense 5G networks has revealed several limitations for vendors and mobile network operators (MNOs), including increased network complexity, high capital expenditure (CapEx), and operational expenditures (OpEx). In addition, energy efficiency has emerged as a critical issue in wireless networks. This is primarily due to the substantial energy consumption of the radio access network (RAN), which involves a dense deployment of base stations (BS) and large-scale multiple-input multiple-output (MIMO) systems at each site. More importantly, the coexistence of legacy and non-cloud-native applications, combined with users’ requirements, has further highlighted the need for more intelligent, flexible, and cost-effective solutions in future 6G systems. Specifically, these challenges have motivated significant research into novel network architectures and communication schemes that aim to: (i) Meet the network performance requirements in terms of coverage, data rates, and latency; (ii) Support the envisioned applications such as integrated sensing and communication (ISAC), artificial intelligence (AI) agent-driven interactions, and cognitive cities; (iii) Enable the deployment of energy-efficient, scalable, and flexible networks; and (iv) Integrate intelligence and data to optimize the network performance and improve user experience with a variety of customized services.
@@
The evolution of the RAN towards an open, disaggregated, and intelligent framework, i.e., Open RAN (ORAN), represents a key step in this direction [10]. Meanwhile, since B5G and 6G systems become increasingly integrated into important application scenarios, such as smart cities or healthcare, any service disruption can lead to serious socio-economic consequences, especially during emergencies or large-scale failures [68]. To extend coverage and enhance overall network resilience, the terrestrial network (TN) can be effectively complemented by a non-terrestrial network (NTN), which provides complementary and redundant communication links to achieve ubiquitous connectivity [18, 4]. Nevertheless, the deployment and operation of NTNs, as well as their interoperability with TN, present several challenges. These include the size, weight, and power (SWaP) constraints, excessive propagation delays, and mobility management issues. These challenges have motivated increasing interest in the integration of ORAN and NTNs. This integration can leverage ORAN’s inherent features of disaggregation, openness, and AI-Native intelligence to enable more efficient network optimization and innovation [35, 103]. However, due to the aforementioned challenges, the adoption of ORAN principles in NTN remains significantly more complicated than in the traditional TN, thus requiring further studies on effective architectural and orchestration solutions.
1-B REVIEW OF EXISTING RELATED LITERATURE
@@
ORAN-based NTN has recently gained significant attention and research interests, with many research papers advancing the idea through investigations on different aspects. It should be noted that the scope of this subsection does not aim to cover the entirety of the existing works in ORAN for NTN, but rather to specifically review the research papers related to state-of-the-art development and deployment solutions of ORAN-based NTN. Therefore, a detailed content comparison between this survey paper and the most relevant and recent research papers is provided in Table 2.
@@
Table 2: Comparison of related works and this paper on ORAN-based NTN (L: Low, M: Medium, H: High).
@@
[24] 2025 Survey M L H H L M Analyzes functional splits and RIC placement for ORAN-enabled satellite-based NTNs, highlighting the implementation constraints and interoperability requirements.
[87] 2024 H M H M L M Provides a detailed overview of the utilization of network slicing, AI/ML techniques, and ORAN to address the challenges of NTN systems in both academia and industry.
[85] 2024 L L M M H M Investigates the existing ORAN testbeds to support UAVs with controlled mobility, and evaluates AERPAW as a reference platform for future 6G NTN experimentation.
[78] 2025 Magazine M L H M L M Proposes an ORAN-based NTN architectures with transparent or regenerative payloads configurations, and outlines the existing challenges in channel, computing and algorithms.
[23] 2025 M M H H M M Proposes the Space-O-RAN framework, integrating satellite-based NTN and TN based on ORAN and AI-driven solutions, to address dynamic connectivity and scalability issues.
[47] 2025 M L M M H H Proposes the Orbital O-RAN (O2-RAN) framework to support multi-layer, multi-segment NTN deployment, and enable intelligent and resilient operations across NTN infrastructures.
[110] 2025 M L L H L L Evaluates different functional split options in ORAN-enabled UAV-based NTNs for URLLC services in terms of computation complexity, latency, and reliability.
[133] 2025 M L H M L M Proposes the SEM-NTN framework, integrating semantic communication, ORAN, and NTNs, for real-time, bandwidth-constrained and AI-Native applications.
[82] 2025 M L M M L M Proposes an AERPAW-based ORAN experimentation platform prototype for wireless UAV research, and analyzes the key challenges and practical solutions for the platform.
[113] 2024 M L L M L L Discusses the options and challenges for deploying ML solutions on ORAN-based TN and NTN scenarios to support centralized and distributed intelligent solutions.
[30] 2024 L L L M H L Discusses the utilization of ORAN architecture and spectrum marketplaces for facilitating spectrum-sharing strategies between NTNs and TN.
[32] 2024 L L M L L M Investigates the ORAN-based NTN architecture in enabling network optimization against various targets, and presents a neutral host 5G network testbed for evaluating the architecture.
[26] 2023 L L M L L M Explores the implementation of ORAN-based NTN architecture and investigates the potential trends, such as radio resource management, to enhance NTN system efficiency.
[39] 2025 White Paper L M H M M L Presents a detailed overview of the integration of ORAN architecture and 3GPP NTN, emphasizing its importance, current status, challenges, and security considerations.
[97] 2025 Book Chapter L L M M H L Summarizes the 3GPP-based NTN specifications, challenges, and various use cases, showing the benefits for realizing the NTN solutions based on ORAN architecture.
@@
Some of these publications provide a detailed overview of the integration of ORAN and NTNs. For instance, the survey paper [24] investigated the architectural and functional split strategies and the RAN intelligent controller (RIC) placement options for ORAN-enabled satellite-based NTN systems, and emphasized the implementation constraints and interoperability considerations. Another in-depth survey paper [87] provided a detailed overview of the utilization of network slicing, AI or machine learning (ML) techniques, and ORAN to address the challenges of NTN systems in both academia and industry. Furthermore, the survey paper [85] discussed the design requirements of ORAN testbeds for controlled experimentation with UAV-based aerial clients, presented the representative experiments from aerial experimentation and research platform for advanced wireless (AERPAW) related to ORAN, and highlighted the AERPAW platform as a potential future ORAN-based UAV testbed.
Furthermore, several magazine papers focused on addressing the overall architecture related to ORAN-based NTNs [78, 23, 47]. Specifically, the authors in [78] proposed an ORAN-empowered NTN architecture with transparent, regenerative-DU, and regenerative payload configurations, and the authors in [23] proposed the Space-ORAN architecture framework by integrating satellite-based NTN and TN based on ORAN principles and AI-driven solutions. In [47], a unified architectural framework named Orbital ORAN (O2-RAN) is proposed to support multi-layer, multi-segment NTN deployment. In addition, other magazine papers have investigated specific aspects of ORAN-enabled NTNs, including ultra reliable low latency communications (URLLC) services [110], semantic communication [133], machine learning solution deployment [113], spectrum sharing [30], and the implementation or testbed designs [82, 32, 26] based on AERPAW. Finally, the white paper [39] analyzed the integration of ORAN architecture and 3GPP-based NTN systems in detail, and the book chapter [97] summarized the relevant characteristics for 3GPP NTN systems and highlighted the benefits of integrating NTN with ORAN architecture. Other surveys or magazine papers with limited relevance to the scope of ORAN for NTN can be found in [123, 84, 117, 115, 100].
@@
1-C CONTRIBUTIONS
In this paper, we provide a comprehensive overview of the ORAN solutions for NTNs and analyze the potential of the AI-Native system in enhancing communication performance. Unlike related studies [24, 87, 85, 78, 23, 47, 110, 30, 133, 113, 82, 32, 26, 39, 97], we analyze the unique DevOps challenges in the NTN lifecycle, propose an AI-Native ORAN-based NTN framework with key technological enablers, present various use cases, and discuss the future directions. The contributions of this paper are summarized as follows:
  • We provide an up-to-date and structured review of the most recent and related studies on the ORAN-based NTNs, highlighting their current progress, contributions, and limitations. In addition, we provide a concise background on ORAN, NTN, and AI-Native for communication, and also present key survey and magazine papers for readers interested in exploring these topics further.
  • We identify and analyze the unique challenges across the DevOps lifecycle of NTNs, including specifications and constraints, robust connectivity, platform control and data analytics, and on-orbit test and upgrade. More importantly, we discuss the corresponding ORAN-based solutions by leveraging its inherent features to address these challenges.
  • We then propose a novel orchestrated AI-Native ORAN-based NTN framework to enable dynamic, scalable, and intelligent network operation for future 6G networks. We also highlight its key enabling technologies, including the dynamic wireless fronthaul (FH) split, multi-tier RIC structure, flexible functional role switch, and cross-domain service management and orchestration (SMO).
@@ 1-C CONTRIBUTIONS
  • We discuss the future research directions on integrating the proposed ORAN-based NTN framework with other technologies, including infrastructure sharing, hierarchical ML-based optimization, edge AI, cross-domain AI, mobility management, and digital twin. These directions will further shape the evolution of AI-driven intelligent wireless networks.
1-D ORGANIZATION
The remainder of this paper is organized as follows. Section 2 provides the detailed background for ORAN, NTN, and AI-Native for communication, and reviews the latest studies related to these topics. Section 3 analyzes the unique challenges of NTN throughout the DevOps lifecycle and discusses corresponding ORAN-based solutions. Section 4 proposes a novel AI-Native ORAN-based NTN framework designed to achieve dynamic and intelligent network configuration, and introduces its key enabling technologies. Section 9 provides three representative use cases empowered by the proposed framework. Section 6 outlines future research directions for integrating the proposed framework with other emerging technologies. Finally, Section 7 concludes the paper. An overview of the paper organization is shown in Fig. 1, and the list of key abbreviations used is provided in Table 1.